(a) Find the derivative of the function 𝑦 = 2π‘₯^2+12/x^2, when π‘₯ = 2.

(b) Let 𝑓(π‘₯) = βˆ’3/π‘₯βˆ’7. Find the inverse of the function.


1
Expert's answer
2022-03-11T12:47:50-0500

Solution (a)

"𝑦 = 2𝑥^2+\frac{12}{x^2}"


"\ud835\udc66 = 2\ud835\udc65^2+12x^{-2}"


"\frac{dy}{dx}= (2)(2)𝑥^{2-1}+(12)(-2)x^{{-2-1}}"


"\frac{dy}{dx}= 4𝑥-24x^{-3}"


at the point "x=2"


"\frac{dy}{dx}= 4(2)-(24)(2)^{-3}"


"\frac{dy}{dx}= 12-\frac{24}{8}"


"\frac{dy}{dx}= 9"


Hence the derivative of "\ud835\udc66 = 2\ud835\udc65^2+12\/x^2", when "\ud835\udc65 = 2" is


"\frac{dy}{dx}= 9"




Solution (b)


Given that "𝑓(𝑥) =y= -\frac{3}{x-7}" .


Interchanging "x" (domain) and "y" (range)


"x= -\frac{3}{y-7}"


Now making "y" the subject of the formula, we have


"x(y-7)=-3"


"y-7=-\frac{3}{x}"


"y=-\frac{3}{x}+7"


Hence the inverse function of the given function "𝑓(𝑥) =y= -\frac{3}{x-7}" is


"𝑓^{-1}(𝑥) =-\frac{3}{x}+7"